The shorter parallel side of a trapezoid is equal to its height, and the longer side is 3/2 times as long. If the area is 80 ft 2, find the length of the shorter side.
Answer by user_dude2008(1862)

A=(h/2)(b1+b2)

80=(x/2)(x+3/2x)

80=5/4x^2

320=5x^2

64=x^2

x=8 or x=-8

Answer: 8 ft